# BayesSDF: Surface-Based Laplacian Uncertainty Estimation for 3D Geometry with Neural Signed Distance Fields

### v2

@@ Abstract
Quantifying uncertainty in neural implicit 3D representations, particularly those utilizing Signed Distance Functions (SDFs), remains a substantial challenge due to computational inefficiencies, scalability issues, and geometric inconsistencies. Existing methods typically neglect direct geometric integration, leading to poorly calibrated uncertainty maps. We introduce BayesSDF, a novel probabilistic framework for uncertainty quantification in neural implicit SDF models, motivated by scientific simulation applications with 3D environments (e.g., forests) such as modeling fluid flow through forests, where precise surface geometry and reliable uncertainty estimates are essential. Unlike radiance-based models such as Neural Radiance Fields (NeRF) or 3D Gaussian splatting, which lack explicit surface formulations, Signed Distance Functions (SDFs) define continuous and differentiable geometry, making them better suited for physical modeling and analysis. BayesSDF leverages a Laplace approximation to quantify local surface instability using Hessian-based metrics, enabling efficient, surfaceaware uncertainty estimation. Our method shows that uncertainty predictions correspond closely with poorly reconstructed geometry, providing actionable confidence measures for downstream use. Extensive evaluations on synthetic and real-world datasets demonstrate that BayesSDF outperforms existing methods in both calibration and geometric consistency, establishing a strong foundation for uncertainty-aware 3D scene reconstruction, simulation, and robotic decision-making.
@@ 1 Introduction
Uncertainty quantification is a fundamental challenge in machine learning, particularly in 3D scene reconstruction, where ensuring reliability in learned representations is crucial for downstream applications such as robotics, medical imaging, and physical simulations. Neural implicit representations have revolutionized 3D geometry learning by encoding scenes as continuous functions rather than discrete point clouds or meshes. Among these, Neural Radiance Fields (NeRF) [5] enable photorealistic novel view synthesis but lack explicit geometric constraints, making them less suitable for precise surface reconstruction. To address this limitation, Neural Implicit Surfaces (NeuS) [7] employ Signed Distance Functions (SDFs), providing a more structured representation of 3D geometry by modeling surfaces as zero-level sets of a continuous field. This geometric formulation allows NeuS to achieve superior surface consistency and geometric regularity, but the challenge remains in effectively quantifying uncertainty in the learned SDF field. In this paper, we introduce BayesSDF, a novel probabilistic framework that enables surface-aligned uncertainty estimation in SDF-based 3D neural representations. Rather than relying on ensembles or pixel-level sampling, BayesSDF utilizes a Laplace approximation over a learned deformation field to compute a Hessian-based metric that captures the local sensitivity of the rendering function. This formulation ensures that uncertainty maps are geometrically coherent. By integrating the geometric advantages of SDF-enabled models with a principled probabilistic framework, BayesSDF establishes an unexplored view for uncertainty-aware 3D scene reconstruction and representation learning. The method enables precise identification of areas with high reconstruction variance, leading to improved surface-aware uncertainty estimates. BayesSDF also improves calibration and geometric consistency over previous methods. Instead of estimating uncertainty through pure pixel-space renderings or depth-based sampling like NeRF, BayesSDF utilizes the continuous SDF field from geometrically adaptable reconstruction models, reducing computational overhead while maintaining accuracy. The proposed method demonstrates superior performance across multiple synthetic and real-world datasets, achieving state-of-the-art uncertainty calibration and reliability in 3D surface modeling. This research introduces significant contributions to the field of uncertainty quantification for 3D implicit neural representations:
2 Contributions
@@
    Direct Geometric Integration: Unlike existing techniques that indirectly approximate uncertainty through pixel-space or depth-based approaches, BayesSDF explicitly integrates the geometric properties inherent in Signed Distance Functions. This direct integration significantly improves the geometric consistency and calibration of uncertainty predictions.
@@
    Efficient Hessian-based Uncertainty Estimation: BayesSDF employs a Hessian-based uncertainty calculation method that efficiently captures local sensitivity in the geometry deformation field. By evaluating how small perturbations in geometric parameters affect rendered color outputs, the method accurately quantifies spatial uncertainty, avoiding the computational expense of traditional ensemble or extensive sampling approaches.
@@
    Scalable and Computationally Efficient Framework: Leveraging a hierarchical hash-based deformation field, BayesSDF maintains high accuracy while reducing computational and memory overhead. The approach integrates seamlessly with existing neural implicit surface frameworks such as NeuS, ensuring a compact yet powerful uncertainty estimation model.
  4. 4.
    Comprehensive Evaluation and Validation: Extensive empirical validation on synthetic and real-world datasets demonstrates BayesSDF’s superiority over existing uncertainty quantification methods. Metrics such as the Area Under Sparsification Error (AUSE) show that BayesSDF achieves comparable or superior calibration performance, especially in complex real-world scenarios.
  5. 5.
    Scientific Motivation for SDF-Based Uncertainty: We motivate our approach by targeting applications in scientific simulation and computational physics, where accurate geometry and interpretable uncertainty are critical. SDF-based models, as demonstrated in DeepSDF [6], provide surface continuity and differentiability that are advantageous for physical modeling. Our framework explicitly supports such use cases by relating uncertainty to local surface instability.
These contributions establish BayesSDF as a robust, geometrically coherent, and computationally efficient uncertainty quantification framework with significant implications for scene reconstruction, physics-based simulations, and autonomous robotic navigation.
3 Related Work
Uncertainty quantification in neural implicit representations is a rapidly evolving field. Traditional deep learning models often produce deterministic outputs without explicit confidence measures, making them unreliable in critical scenarios. In the context of 3D learning, uncertainty quantification plays a crucial role in assessing reconstruction accuracy, identifying ambiguous regions, and guiding active learning. Below, we review the most relevant areas of uncertainty quantification in 3D learning, including ensemble-based methods, neural radiance fields (NeRF), neural implicit surfaces (NeuS), signed distance functions (SDFs), Bayesian uncertainty estimation techniques, among others. We then highlight how BayesSDF significantly advances the field by integrating uncertainty directly into SDF-based neural implicit models, enabling computationally efficient and geometrically consistent uncertainty estimation.
@@ 3.1 Uncertainty Estimation in 3D Reconstruction Methods
Uncertainty estimation plays a crucial role in 3D reconstruction, where accurate surface modeling is essential for downstream applications. Various methods exist for reconstructing 3D surfaces, and each presents unique challenges for uncertainty quantification.
@@ 3.1.1 Neural Representations and SDF-Based Methods
Neural implicit representations, such as NeRF (Neural Radiance Fields) and NeuS (Neural Implicit Surfaces), have demonstrated state-of-the-art performance in view synthesis and geometric reconstruction. However, their uncertainty estimation techniques differ significantly.
  • NeRF-based methods estimate uncertainty per-ray by analyzing the variance in color and density outputs across multiple samples along a viewing ray. These approaches are effective for rendering applications but fail to capture geometric uncertainty directly.
  • NeuS-based methods instead represent surfaces using a SDF, which provides a continuous and differentiable description of 3D geometry. However, existing NeRF uncertainty estimation techniques do not translate well to NeuS because they rely on ray-space density variations rather than surface-level geometric uncertainty.
3.2 Uncertainty Estimation in Ensemble-Based Learning
Ensemble-based uncertainty estimation has emerged as a powerful technique in deep learning, providing a practical means of approximating epistemic uncertainty. In the context of 3D learning and reconstruction methods, ensemble methods are particularly advantageous as they offer robust uncertainty quantification without requiring complex probabilistic modeling or extensive sampling. While ensemble methods are widely used in 2D vision tasks (e.g. image classification and segmentation), applying them effectively in 3D neural representations presents unique challenges:
  • Computational Cost: Training multiple neural implicit models is memory- and computeintensive, particularly for high-resolution 3D reconstructions. Traditional ensemble methods require storing and evaluating multiple models, leading to scalability issues.
  • Surface Consistency: Unlike 2D image-based tasks, uncertainty in 3D learning must be estimated consistently across a continuous surface rather than at discrete pixel locations
  • Ray vs. Volume-Based Estimation: Many existing ensemble-based uncertainty estimation techniques in NeRF-like models operate strictly along rays. This results in uncertainty that does not align well with geometric structures, making it less interpretable.
While leveraging the strengths of ensemble-based learning and addressing key limitations in 3D implicit representations, BayesSDF establishes a new paradigm for uncertainty quantification in neural implicit fields.
@@ 3.3 Bayesian Inference for Ray-Based Sampling
BayesRays [2] is a recent approach for uncertainty quantification in neural radiance fields. It modifies NeRF by introducing a Bayesian framework that estimates uncertainty in the color and density outputs of the model. While BayesRays is effective for quantifying NeRF uncertainty, it remains tied to volumetric density representations and does not explicitly model surface uncertainty. Additionally, BayesRays relies on ray-space sparsification techniques, which can lead to misalignment between estimated uncertainty and actual geometric uncertainty. Unlike BayesRays, BayesSDF integrates reconstruction methods that inherently reveal volumetric qualities, ensuring direct alignment between uncertainty estimates and the scene geometry. This approach places less emphasis on rendering estimations, making it inherently more interpretable for geometric reconstruction tasks by exploiting surface variability rather than depth-based uncertainty.
@@
NeRF Volume Density None High Low
NeuS SDF None High High
BayesRays Volume Density Ensemble Variance Low Moderate
BayesSDF Volume Density Ensemble Variance Low Very High
Table 1: Comparison of existing 3D reconstruction methods, evaluating their representation, uncertainty estimation, computational cost, and geometric consistency.
@@ 4 Background
Quantifying uncertainty in 3D implicit representations is essential for improving the reliability and interpretability of learned geometric models. This section reviews the foundational components upon which BayesSDF builds: Signed Distance Functions (SDFs), Neural Radiance Fields (NeRF), Neural Implicit Surfaces (NeuS), ensemble-based uncertainty methods, and Bayesian estimation frameworks such as BayesRays.
4.1 Signed Distance Functions (SDFs)
@@
$f(\mathbf{x})=\begin{cases}\phantom{-}d,&\text{if }\mathbf{x}\text{ is outsidethe surface}\\ \phantom{-}0,&\text{if }\mathbf{x}\text{ lies on the surface}\\ -d,&\text{if }\mathbf{x}\text{ is inside the surface}\end{cases}$
This representation allows for accurate computation of surface normals and curvature, and forms the basis of many modern 3D reconstruction methods [6]. Despite their geometric advantages, SDFs have not been fully explored for probabilistic modeling or uncertainty estimation in learned 3D representations.
4.2 Neural Radiance Fields (NeRF)
NeRF models learn a volumetric function that maps 3D positions and viewing directions to emitted color and volume density [5]. The function $F_{\theta}:(\mathbf{x},\mathbf{d})\rightarrow(\sigma,\mathbf{c})$ enables photorealistic novel view synthesis by rendering images through volume integration:
@@ 4.2 Neural Radiance Fields (NeRF)
[…]
4.3 Neural Implicit Surfaces (NeuS)
NeuS addresses the geometric limitations of NeRF by replacing volumetric density fields with SDFs that model surfaces as zero-level sets [7]. The rendering is adapted by transforming SDF values into opacity via a sigmoid function:
@@ 4.3 Neural Implicit Surfaces (NeuS)
NeuS enables precise geometric modeling, but lacks inherent mechanisms for uncertainty quantification.
@@ 4.4 Uncertainty Estimation via Ensembles
This variance captures model disagreement, which is indicative of uncertainty [4]. While ensemble methods are effective across 2D and 3D vision tasks [1], they are resource-intensive and can be misaligned with underlying geometry in volumetric or ray-based models.
4.5 Bayesian Uncertainty in NeRF: BayesRays
BayesRays proposes a Bayesian approach for NeRF by learning a density-aware variance model and propagating it along rays [2]. The estimated uncertainty is given by:
$U(\mathbf{x})=\frac{1}{K}\sum_{k=1}^{K}\sigma_{k}(\mathbf{x})\cdot T_{k}(\mathbf{x})$
Color and density are modeled as Gaussian random variables, and uncertainty in the rendered image is computed by propagating these distributions through the rendering equation:
$\text{Var}[\hat{C}(\mathbf{r})]=\sum_{i=1}^{N}T_{i}^{2}\left(1-\exp(-\sigma_{i}\delta_{i})\right)^{2}\text{Var}[\mathbf{c}_{i}]$
BayesRays achieves efficient uncertainty estimation without ensembles, but its output lacks surface-level geometric consistency. Integrating such uncertainty with geometric fields remains a core challenge in the field [3].
@@ 5 Method
In this section, we detail our proposed BayesSDF method, which leverages a NeuS-like scene representation and adopts an uncertainty estimation framework through a deformation field and a Hessian-based metric. We begin by describing the theoretical underpinnings and motivation, then present how we construct and query the deformation field, calculate the Hessian, and finally approximate local uncertainty.
@@ 5.1 Theory
Our goal is to model an associated uncertainty measure, derived from the local sensitivity of the rendering function to small perturbations in the underlying deformation parameters. Let $\mathbf{x}\in\mathbb{R}^{3}$ denote a 3D point in scene space, and let $\mathbf{c}(\mathbf{x})\in\mathbb{R}^{3}$ be the color (in Red Green Blue (RGB)) rendered by the neural field. Rather than relying purely on a deterministic mapping $\mathbf{x}\mapsto\mathbf{c}(\mathbf{x})$, we attach a set of local deformation parameters $\mathbf{d}(\mathbf{x})$ that influence the geometry (and thus the color) through a learned distortion or deformation field. Formally, we assume $\mathbf{c}(\mathbf{x})\;=\;\mathcal{R}\bigl{(}\mathbf{x},\mathbf{d}(\mathbf{x});\Theta\bigr{)}$ where $\mathcal{R}$ is the rendering operator that depends on both the scene geometry (parameterized by a neural SDF model and any additional shape parameters $\Theta)$ and the local deformation $\mathbf{d}(\mathbf{x})$. To estimate uncertainty, we measure how sensitively $\mathbf{c}(\mathbf{x})$ changes with respect to changes in $\mathbf{d}(\mathbf{x})$. Intuitively, if small variations in $\mathbf{d}(\mathbf{x})$ induce large changes in color, we interpret that region as having higher uncertainty in the final scene reconstruction.

### v3

@@ Abstract
Accurate surface estimation is critical for downstream tasks in scientific simulation, and quantifying uncertainty in implicit neural 3D representations still remains a substantial challenge due to computational inefficiencies, scalability issues, and geometric inconsistencies. However, current neural implicit surface models do not offer a principled way to quantify uncertainty, limiting their reliability in real-world applications. Inspired by recent probabilistic rendering approaches, we introduce BayesSDF, a novel, probabilistic framework for uncertainty estimation in neural implicit 3D representations. Unlike radiance-based models such as Neural Radiance Fields (NeRF) or 3D Gaussian Splatting, Signed Distance Functions (SDFs) provide continuous, differentiable surface representations, making them especially well-suited for uncertainty-aware modeling. BayesSDF applies a Laplace approximation over SDF weights and derives Hessian-based metrics to estimate local geometric instability. We empirically demonstrate that these uncertainty estimates correlate strongly with surface reconstruction error across both synthetic and real-world benchmarks. By enabling surface-aware uncertainty quantification, BayesSDF lays the groundwork for more robust, interpretable, and actionable 3D perception systems.
@@ 1 Introduction
Scientific applications such as fluid simulation, medical reconstruction, and robotic perception require not only accurate surface reconstructions but also reliable uncertainty estimates [1] to support robust decision-making (e.g. aleatoric vs epistemic [14]). Recent advances in multi-view 3D reconstruction using neural implicit representations [20] have dramatically improved surface quality, yet these methods are rarely adopted in the real world [19] due to their inability to model surface uncertainty. Neural implicit representations model geometry as continuous functions [20, 3, 4], enabling high-resolution reconstructions without relying on dense voxels or explicit meshes. While these methods, like Neural Radiance Fields [20], are effective at view synthesis, they lack surface-level structure and provide no mechanism for quantifying geometric uncertainty. Models like NeuS [47] (and dynamic topological models like Nerfies [30], HyperNeRF [31], CamP [park2022campnerf], etc.) address this limitation by using Signed Distance Functions (SDFs), which define surfaces as zero-level sets of continuous scalar fields. This provides strong geometric structure and differentiability, enabling more consistent and physically meaningful reconstructions. However, these models remain deterministic and offer no mechanism for estimating uncertainty over the learned surface SDF fields. To fill this gap, we introduce BayesSDF, a probabilistic framework for estimating surface-aligned uncertainty in neural implicit models that use SDFs. Our work is inspired by BayesRays [10], which introduced uncertainty modeling for volumetric radiance fields via Bayesian rendering. However, BayesSDF departs fundamentally from BayesRays by focusing on explicit surface geometry. Instead of modeling uncertainty through pixel-space or volumetric sampling, BayesSDF leverages a Laplace approximation [9] over SDF weights to quantify local surface sensitivity using second-order derivatives. This results in uncertainty maps that are geometrically coherent and directly tied to surface stability. By combining the structure of SDF-based models with a tractable probabilistic approximation, BayesSDF produces uncertainty estimates that strongly correlate with reconstruction error, offering more interpretable and actionable outputs. These estimates help identify ambiguous or ill-constrained geometry, which is critical for downstream tasks that require reliable spatial reasoning [50]. Unlike uncertainty models based on pixel-space rendering or volumetric sampling, BayesSDF provides surface-aware estimates rooted in the geometry itself. This enables more efficient and scalable modeling without sacrificing spatial precision. To summarize, our key contributions are as follows:
  • Surface-Aware Uncertainty via SDF Geometry: Prior methods often model uncertainty indirectly in pixel or depth space, which can misalign with true geometric errors. BayesSDF instead leverages the analytic gradients of Signed Distance Functions to estimate uncertainty directly on the surface. This direct integration enables uncertainty estimates that are physically meaningful and geometrically consistent.
  • Uncertainty Estimation through Local Surface Sensitivity: BayesSDF uses a Laplace approximation and Hessian-based calculations over the neural SDF weights, capturing how local geometric perturbations affect surface stability in the geometry’s deformation field. By evaluating how small perturbations in geometric parameters affect rendered color outputs, the method accurately quantifies spatial uncertainty, avoiding the computational expense of traditional ensemble [17] or extensive sampling approaches.
  • Scalable Deformation Modeling with Hash Encoding: To model local geometric variation, BayesSDF introduces a hierarchical multi-resolution hash-encoded deformation field. Compared to grid-based representations used in prior probabilistic methods like BayesRays [10], our approach supports higher spatial resolution with significantly reduced memory overhead. This design makes BayesSDF scalable to complex scenes while remaining compatible with frameworks like NeuS [47].
Motivated by computational physics and scientific simulation applications with 3D environments such as modeling fluid flow through forests, where precise surface geometry and reliable uncertainty estimates are essential. SDF-based models, as demonstrated in DeepSDF [29], provide surface continuity and differentiability that are advantageous for physical modeling. The proposed method demonstrates superior performance across multiple synthetic and real-world datasets, achieving state-of-the-art uncertainty calibration and reliability in 3D surface modeling. Our framework explicitly supports such use cases by relating uncertainty to local surface instability.
2 Related Work
Uncertainty estimation is essential in 3D learning systems, especially for safety-critical and scientific applications, where downstream tasks depend on the reliability of reconstructed geometry. Traditional neural implicit methods such as Neural Radiance Fields (NeRF) [20, 19] and Signed Distance Function (SDF)-based models like NeuS [47] have achieved high-fidelity reconstruction [59, 44] and view synthesis, but largely ignore uncertainty modeling. This omission limits their usefulness in tasks requiring interpretable or risk-aware decision-making. In this section, we contextualize BayesSDF, within three major threads of prior work:
@@ 2 Related Work
    Why SDF-based models [29, 47] offer superior structure for geometric uncertainty estimation compared to NeRF-style density fields. [20, 19, 7]
@@
    How uncertainty has been previously modeled in neural 3D systems, including ensemble methods and Bayesian inference.
@@
    Why BayesSDF provides a principled and non-trivial improvement over prior Bayesian extensions like BayesRays [10], particularly in aligning uncertainty with surface structure.
@@ 2.1 Uncertainty Estimation in 3D Reconstruction Methods
3D reconstruction methods vary widely in how they represent geometry, which directly impacts how uncertainty can be quantified. NeRF-style models [20, 3, 4] rely on volume densities sampled along camera rays. These are well-suited for view synthesis but decouple uncertainty from physical surfaces. In contrast, SDF-based models [29, 39, 36, abou2022freenerf, 40, 12] define surfaces as the zero level-set of continuous scalar fields, offering geometric interpretability and a natural differentiable structure [28, 33, 54], which are features critical for scientific computing tasks involving simulations, collisions, or boundary conditions. BayesSDF builds on this structure by estimating uncertainty not in ray-space but in the SDF’s parameter space. This enables surface-aware uncertainty aligned with real geometric features rather than visual artifacts.
@@ 2.1.1 Neural Representations and SDF-Based Methods
Neural implicit representations, such as NeRF [20] and NeuS [47], have demonstrated state-of-the-art performance in view synthesis and geometric reconstruction. However, their uncertainty estimation techniques differ significantly.
  • NeRF-based methods [20, 37, 38] estimate the uncertainty per ray by analyzing the variance in color and density outputs across multiple samples along a viewing ray. These approaches are effective for rendering applications, but they do not capture geometric uncertainty directly [2].
  • NeuS-based methods [47] instead represent surfaces using an SDF, which provides a continuous and differentiable description of 3D geometry. [41, 13] However, existing NeRF [20] uncertainty estimation techniques do not translate well to NeuS [47] because they rely on ray-space density variations rather than surface-level geometric uncertainty.
2.2 Ensemble-Based Methods and Their Limitations
Ensemble-based methods [17, 1, 14] are a standard approach to approximate epistemic uncertainty by training and sampling multiple networks. In the context of reconstruction methods, ensemble methods are particularly advantageous as they offer robust uncertainty quantification without requiring complex probabilistic modeling or extensive sampling. While ensemble methods are widely used in 2D vision tasks (e.g. image classification and segmentation), they introduce several limitations in the 3D domain:
  • Scalability: Training multiple neural implicit models is memory and compute intensive, particularly for high-resolution 3D reconstructions. [23, 8, 55, 45] Traditional ensemble methods require storing and evaluating multiple models, leading to scalability issues [48].
  • Surface Consistency: Unlike 2D image-based tasks [21], uncertainty in 3D learning must be estimated consistently across a continuous surface rather than at discrete pixel locations. [61, 58]
  • Ray vs. Volume-Based Estimation: Many existing ensemble-based uncertainty estimation techniques in NeRF-like [20] models operate strictly along rays. This results in uncertainty that does not align well with geometric structures, making it less interpretable.
These challenges motivate more principled Bayesian approximations within neural 3D representations leveraging the strengths of ensemble-based learning, particularly those offering geometric continuity like SDFs [53].
@@ 2.3 Bayesian Inference for Ray-Based Sampling
BayesRays [10] introduced uncertainty estimation in NeRF [20] by modeling pixel-space color and density distributions using a Bayesian ensemble and ray sparsification. While BayesRays [10] is effective for capturing radiance uncertainty, it remains tied to volumetric density representations and does not explicitly model surface uncertainty [37, 32, 6, 60, 34, 11]. It relies on ray-space sparsification techniques [30, 31, 25, 46], which can lead to misalignment between estimated uncertainty and actual geometric uncertainty [43]. Its volumetric density representation lacks an explicit notion of surface, making it harder to isolate geometric ambiguity. BayesSDF addresses these issues by modeling the sensitivity of the surface geometry itself. Rather than relying on variation in rendered pixel values, BayesSDF quantifies how local perturbations in the SDF field affect surface stability using a Laplace approximation [9, 39, 36, 26, 52, 35]. This results in uncertainty maps that are tightly coupled to the surface [29, 24], making them more interpretable and physically meaningful. Notably, directly applying BayesRays [10] to SDF-based models like NeuS [47] is non-trivial. BayesRays [10] relies on sampling distributions along rays [29, 27], which do not generalize to SDF’s zero-crossing surface logic. In contrast, BayesSDF incorporates the analytic structure of SDFs via second-order derivatives [9, 39, 24], enabling uncertainty estimates that are both tractable and geometrically consistent.
@@
NeRF [20] Volume Density None High Low
NeuS [47] SDF None High High
BayesRays [10] Volume Density Ray-Space Variance Moderate Moderate
BayesSDF SDF Hessian-Based Laplace Approx. Low Very High
Table 1: Comparison of existing 3D reconstruction methods based on geometry representation, uncertainty modeling, computational cost, and surface-level interpretability.
@@ 3 Background
Quantifying uncertainty in 3D implicit representations is essential to improve the reliability and interpretability of learned geometric models. This section reviews the foundational components upon which BayesSDF builds upon.
3.1 Neural Implicit Surfaces (NeuS)
@@
This representation allows for accurate computation of surface normals and curvature, and forms the basis of many modern 3D reconstruction methods [29, 5, 12, 33, 28, 54, 26]. Despite their geometric advantages, SDFs have not been fully explored for probabilistic modeling or uncertainty estimation in learned 3D representations [18, 11, 53]. NeuS [47] addresses the geometric limitations of NeRF [20] by replacing volumetric density fields with SDFs that model surfaces as zero-level sets. The rendering is adapted by transforming SDF values into opacity via a sigmoid function:
@@
NeuS [47] enables precise geometric modeling, but lacks inherent mechanisms for uncertainty quantification.
@@ 3.2 Uncertainty Estimation via Ensembles
This variance captures model disagreement, which is indicative of uncertainty [17, 14, 2, 57]. While ensemble methods are effective across 2D and 3D vision tasks [1], they are resource-intensive and can be misaligned with underlying geometry in volumetric or ray-based models [61, 56].
@@ 4 Method
In this section, we detail our proposed BayesSDF method, which leverages a NeuS [47] scene representation and adopts an uncertainty estimation framework through a Hessian-based deformation field. We begin by describing the theoretical underpinnings, then present how we construct and query the deformation field, calculate the Hessian, and finally approximate local uncertainty.
@@ 4.1 Theory
Our goal is to model an associated uncertainty measure derived from the local sensitivity of the rendering function to small perturbations in the underlying deformation parameters. Let $\mathbf{x}\in\mathbb{R}^{3}$ denote a 3D point in the scene space, and let $\mathbf{c}(\mathbf{x})\in\mathbb{R}^{3}$ be the color (in red-green-blue (RGB)) rendered by the neural field. Rather than relying solely on a deterministic mapping $\mathbf{x}\mapsto\mathbf{c}(\mathbf{x})$, we attach a set of local deformation parameters $\mathbf{d}(\mathbf{x})$ that influence the geometry (and thus the color) through a learned distortion or deformation field [32, 30]. Formally, we assume $\mathbf{c}(\mathbf{x})\;=\;\mathcal{R}\bigl{(}\mathbf{x},\mathbf{d}(\mathbf{x});\Theta\bigr{)}$ where $\mathcal{R}$ is the rendering operator that depends on both the scene geometry (parameterized by a neural SDF model and any additional shape parameters $\Theta)$ and the local deformation $\mathbf{d}(\mathbf{x})$. To estimate uncertainty, we measure how sensitively $\mathbf{c}(\mathbf{x})$ changes with respect to changes in $\mathbf{d}(\mathbf{x})$ [14, 1]. Intuitively, if small variations in $\mathbf{d}(\mathbf{x})$ induce large changes in color, we interpret that region as having higher uncertainty in the final scene reconstruction [9, 35].